Simplify the expression 12x+10 over 2: Simplify the expression 12x+10 over 2
Answer by jojo14344(883)

= (12x+10)/2 becomes,
(12x/2) + (10/2) = 0 ---------> working eqn
further simplifying,
12x/2 + 5 = 0
12x/2=-5
-10=12x
x= (-5/6) --------------> final answer
To check, go back working eqn & substitute the value of "x",
((12)(-5/6))/2 + 10/2 = 0
(-60/6)/2 + 10/2 = 0
 cross(-10/2) + cross(10/2) = 0
0=0
Thank you,
Jojo
Question 149590: Simplify the expression 12x+10 over 2: Simplify the expression 12x+10 over 2
Answer by jim_thompson5910(9376)

(12x+10)/2 Start with the given expression.


(2(6x+5))/2 Factor out the GCF 2 from 12x+10 to get 2(6x+5)


(highlight(2)(6x+5))/highlight(2) Highlight the common terms.



(cross(2)(6x+5))/cross(2) Cancel out the common terms.


6x+5 Simplify.



So (12x+10)/2 simplifies to 6x+5.


In other words, (12x+10)/2=6x+5